OpenStudy (anonymous):

Advance algebra financial applications HELP PLEASE! Yon made contributions to a Traditional IRA over the course of 15 working years. Her contributions averaged $4,000 annually. Yon was in the 29% tax bracket during her working years. The average annual rate of return on the account was 5%. Upon retirement, Yon stopped working and making IRA contributions. Instead, she started living on withdrawals from the retirement account. At this point, Yon dropped into the 15% tax bracket. Factoring in taxes, what is the effective value of Yon’s Traditional IRA at retirement? Assume annual compounding.

OpenStudy (anonymous):

Answer choices $74,230.25 $25,031.13 $73,367.11 $141,367.11

OpenStudy (anonymous):

@ganeshie8 can you walk me through this please?

ganeshie8 (ganeshie8):

okay just gimme 5 minutes brb

OpenStudy (anonymous):

No problem I know im going to use FVOA @ganeshie8

ganeshie8 (ganeshie8):

Yes, use FVOA and find out the accumulated sum first.

ganeshie8 (ganeshie8):

\[ FVOA = C \bullet \frac{(1+i)^{nt}-1}{i} \]

ganeshie8 (ganeshie8):

\(C\) = periodic cash payment = $4,000 \(i\) = periodic interest rate = 0.05 \(n\) = 1 \(t\) = 15

ganeshie8 (ganeshie8):

plug and chug

OpenStudy (anonymous):

the sum is 86,314.25

ganeshie8 (ganeshie8):

yes, you need to cut out the taxes from that

ganeshie8 (ganeshie8):

whats the tax percent ?

OpenStudy (anonymous):

15% so .15*86,314.25= 12,947.14

ganeshie8 (ganeshie8):

yes, u need to give it off to government as tax

ganeshie8 (ganeshie8):

so effective money u wil be having when u retire = ?

OpenStudy (anonymous):

73,367.11

ganeshie8 (ganeshie8):

Perfect !

ganeshie8 (ganeshie8):

You could have gotten that by simply taking 85% of the accumulated money

ganeshie8 (ganeshie8):

instead of calucatign 15% and subtracting..

OpenStudy (anonymous):

Yayy thank you for your help i did another problem like this the same way but got it wrong i dont understand how if it's set up the same way could you take a look at it maybe i used the wrong formula

ganeshie8 (ganeshie8):

15% tax takes away, that means u wil be getting 85% so .85*86,314.25= 73,367.11

ganeshie8 (ganeshie8):

sure, post it :)